How much of a .5M calcium hydroxide is needed to neutralize 250.0ml of a .5M phosphoric acid
1
Expert's answer
2020-06-04T10:30:53-0400

3 Ca(OH)2 + 2 H3PO4 → Ca3(PO4)2 + 6 H2O


0.2500 L * 0.5 M = 0.125 mol of phosphoric acid

the ratio is 3 : 2 so

n (Ca(OH)2) = 3/2 * 0.125 mol = 0.1875 mol

0.1875 mol / 0.5 M = 0,375 L or 375 mL